Out of 100 People 65 are tea drinker while 45 are coffee drinker and 30 are both find the probability that
1. There is either tea or coffee drinker
2. Neither Tea or coffee drinker
1
Expert's answer
2019-10-16T10:49:00-0400

By the inclusion–exclusion principle, we obtain that "65+45-30=80" people drink either coffee or tea. So "100-80=20" people drink neither coffee nor tea.

So a)probability that there is either coffee or tea drinker is "\frac{80}{100}=\frac{4}{5}" and b)probability that there is neither coffee nor tea drinker is "\frac{20}{100}=\frac{1}{5}"

Answer: a)"\frac{4}{5}" b)"\frac{1}{5}"